To open a clothing repair shop, you need to carry out the following expenses: purchase of raw materials and materials - 50 thousand UAH. per year, the cost of equipment - 200 thousand UAH. (service life - 10 years), rental of premises - 3 thousand UAH. monthly, salary of employees - UAH 8.4 thousand. monthly, electricity - 18 thousand UAH. in year.

For the amount of funds invested in production, it would be possible to receive an income of 7% per annum, as well as receive wages in a similar production - 30 thousand UAH. in year.

Calculate accounting, economic and normal profit if the total income of the enterprise is 550 thousand UAH. in year


1
Expert's answer
2021-02-28T11:37:27-0500

Accounting profit is:

AP = 550 - (50 + 200 + (3 + 8.4)×12 + 18) = 145.2 UAH.

Economic profit is:

EP = 145.2 - (404.8×0.07 + 30) = 86.86 UAH.

Normal profit is: NP = 0.